how do I solve the following?
u(square root of (75uy^2)) + 8y (suare root of (48 u^3))

Answer by amoresroy(361)

u(square root of (75uy^2)) + 8y (suare root of (48 u^3))
= u (75uy^2)^1/2 + 8y (48u^3)^1/2
= u (5*5*3*u*y*y)^1/2 + 8y (4*4*3*u*u*u)^1/2
= u*5*y(3u)^1/2 + 8y*4*u(3u)^1/2
= 5uy (3u)^1/2 + 32uy (3u)^1/2
= 37uy(3u)^1/2